# Streaming Federated Learning with Markovian Data

### v1

@@ Appendix F Local SGD with Momentum
Hence, by using Young inequality:
@@
Here in the last step, we use triangle inequality and the sample-wise smoothness (Assumption 4.2).
@@
Using again Young inequality we have:
@@
We start by using Young inequality $\left(a+b\right)^{2}\leq\left(1+\frac{1}{K-1}\right)a^{2}+Kb^{2}$. Then, for any $10\leq k\leq K-$ 1 we have:
$\begin{split}\mathbb{E}\left[\left\lVert w_{t}^{(m,k)}-w_{t}\right\rVert^{2}\right]\leq&\quad\left(1+\frac{1}{K-1}\right)\mathbb{E}\left[\left\lVert w_{t}^{(m,k-1)}-w_{t}\right\rVert^{2}\right]+K\mathbb{E}\left[\left\lVert w_{t}^{(m,k-1)}-w_{t}^{(m,k)}\right\rVert^{2}\right]\\ =&\quad\left(1+\frac{1}{K-1}\right)\mathbb{E}\left[\left\lVert w_{t}^{(m,k-1)}-w_{t}\right\rVert^{2}\right]\\ &+K\eta^{2}\mathbb{E}\left[\left\lVert\beta\nabla f_{m}\left(w_{t}^{(m,k-1)};x_{t}^{(m,k-1)}\right)+\left(1-\beta\right)v_{t}\right\rVert^{2}\right].\\ \end{split}$ (114)
@@
Hence, by using triangle inequality we have:
@@
Here we use the sample-wise smoothness and the uniform bound on the gradient noise in the last step (Assumptions 4.24.3).
@@
With the condition on $\eta$, we have that $4\eta^{2}\beta^{2}L^{2}K\leq\frac{1}{K}<\frac{1}{K-1}$. Hence by using the geometric sum formula, we have:
@@
In the proof of Local SGD we have to use the heterogeneity assumption to bound the term $\frac{1}{M}\sum_{m}\mathbb{E}\left[\left\lVert\nabla F_{m}\left(w_{t}\right)\right\rVert^{2}\right]$. The use of momentum allows us to build a recursive bound for this term by the following lemma.
Lemma F.4.
@@
By using Young inequality and the smoothness of $F_{m}$, we have that:
@@
$\begin{split}\eta KL\leq&\left(\frac{1}{2\beta}\wedge\frac{1}{\sqrt{120MK\beta^{2}}}\wedge\sqrt{\frac{L\Delta_{0}}{360\beta^{3}TG_{0}}}\wedge\frac{1}{225}\frac{1}{\left(1-\beta\right)}\wedge\frac{1}{525}\frac{1}{\beta\gamma LT}\right),\\ \gamma\leq&\frac{1}{\sqrt{60}}\frac{\beta}{L}.\end{split}$ (123)
@@
$\begin{split}\frac{1}{T}\sum_{t=0}^{T-1}\Xi_{t}\leq\frac{7}{2}\frac{L\Delta_{0}}{\beta T}+\frac{17}{2}\frac{\tau\sigma^{2}}{MK}+\frac{1}{4}\sum_{t=0}^{T-1}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+\frac{15}{2}\frac{\sigma^{2}}{T}\sum_{t=0}^{T-1}\left\lVert QP^{Kt}-\pi\right\rVert_{TV}.\end{split}$ (124)
@@
By Lemma F.4, we have that:
@@
Plugging (125) back into Lemma F.3 we have:
@@
By condition on the step size, we have that $\mathcal{B}\leq\frac{2}{15}$. Therefore, plugging (126) back into Lemma F.2, by noticing that $\mathbb{E}\left[\left\lVert\nabla F\left(w_{-1}\right)\right\rVert^{2}\right]=0$, gives:
@@
$\begin{split}120(\eta KL\beta)^{2}\sigma^{2}\leq&\frac{\sigma^{2}}{MK}\leq\frac{\tau\sigma^{2}}{MK}\quad\textrm{(since $\tau$ is always bigger than or equal to $1$)},\\ 360\left(\eta KL\beta\right)^{2}G_{0}\leq&\frac{L\Delta_{0}}{\beta T}.\end{split}$ (129)
Hence, by plugging (129)) and Lemma C.1 back into (128) we have:
$\begin{split}\frac{1}{T}\sum_{t=0}^{T-1}\Xi_{t}\leq\frac{5}{4}\frac{\Xi_{-1}}{\beta T}+\frac{L\Delta_{0}}{\beta T}+\frac{16\tau\sigma^{2}}{MK}+\frac{1}{4}\sum_{t=0}^{T-1}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+\frac{15}{2}\frac{\sigma^{2}}{T}\sum_{t=0}^{T-1}\left\lVert QP^{Kt}-\pi\right\rVert_{TV}.\end{split}$ (130)
Finally noticing that if we initialize $v_{0}=0$, $\Xi_{-1}=\mathbb{E}\left[\left\lVert\nabla F\left(w_{0}\right)\right\rVert^{2}\right]\leq 2L\Delta_{0}$ by the smoothness of $F$. Hence we prove the lemma. ∎
@@
Theorem F.6 (Theorem 4.9 in the main paper).
@@
$\begin{split}\eta KL&\leq\mathcal{O}\left(\frac{1}{\beta}\wedge\frac{1}{\sqrt{MK\beta^{2}}}\wedge\sqrt{\frac{L\Delta_{0}}{\beta^{3}TG_{0}}}\wedge\frac{1}{\left(1-\beta\right)}\wedge\frac{1}{\beta\gamma LT}\right),\\ \gamma&\leq\mathcal{O}\left(\frac{\beta}{L}\right),\end{split}$ (131)
@@
$\begin{split}\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\mathcal{O}&\left(\frac{L\Delta_{0}}{\beta T}+\frac{\tau\sigma^{2}}{MK}\right.\left.+\frac{\sigma^{2}}{T}\sum_{t=0}^{T-1}\left\lVert QP^{Kt}-\pi\right\rVert_{TV}\right).\end{split}$ (132)
@@
By rearranging the terms in Lemma F.1:
@@
Using the condition on $\gamma$ and replacing $\frac{1}{T}\sum_{t=0}^{T-1}\Xi_{t}$ by Lemma F.5 proves the theorem. ∎
From Theorem F.6, we can derive directly the communication and sample complexity of Local SGD-M.
Corollary F.7 (Corollary 4.10 in the main paper).
Under the conditions of Theorem F.6, the required number of local steps and communication rounds for Local SGD-M to achieve $\mathbb{E}[\|\nabla F(\hat{w}_{T})\|^{2}]\leq\epsilon^{2}$ are:
  1. (i)
    If the system-level Markov chain $X$ is stationary ($Q=\pi$),
    $K\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),\leavevmode\nobreak\ T\geq\mathcal{O}\left(\frac{L\Delta_{0}}{\beta\epsilon^{2}}\right).$ (134)
  2. (ii)
    If the system-level Markov chain $X$ is non-stationary ($Q\neq\pi$), with $\epsilon\leq\mathcal{O}\left(\sqrt{\frac{\sigma^{2}}{M}}\right)$,
    $K\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),\leavevmode\nobreak\ T\geq\mathcal{O}\left(\frac{\frac{L\Delta_{0}}{\beta}\vee\sigma^{2}}{\epsilon^{2}}\right).$ (135)
@@
The proof is a direct consequence of Theorem F.6 and follows exactly the same steps in the proof of Corollary D.3. ∎
Appendix G Additional experiments
The following experiments investigate additional aspects of the scenario described in Sec. 5.
Impact of Heterogeneity: We complete the experiments in Sec. 5. In Figures 34, we plot the gradient norm trajectory $\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}$ for Minibatch SGD, Local SGD, and Local SGD-M, with $\gamma=0.01,\eta=0.001$, $\beta=0.1$, the number of clients $M=\in\left\{10,100\right\}$, the number of local steps $K\in\{10,100,1000\}$, and the mixing time for every client $\tau_{m}\in\left\{10,1000\right\}$. As predicted by our analysis, the performance Local SGD stalls when the number of local updates increases, while Minibatch SGD and Local SGD-M are more robust to the heterogeneity. In particular, performance is improved as the number of samples $K$ in each communication round is increased.
Impact of Number of Clients $M$: We complete the experiments in Sec. 5. Figure 5 plots the gradient norm trajectory $\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}$ for Minibatch SGD, Local SGD, and Local SGD-M, and varying numbers of clients. The algorithm parameters are given by: $\gamma=0.01,\eta=0.001$, $\beta=0.1$, $\lambda=0.01$, and the number of local steps $K=100$. Observe that the gradient norm for all three algorithms decreases when the number of clients increases, indicating that collaboration is beneficial even with heterogeneous Markovian data.
Effect of mixing time: In Figure 6, we plot the gradient norm of the last iterate of Minibatch SGD, Local SGD, and Local SGD-M, with $\gamma=0.01,\eta=0.001$, $\beta=0.1$, $\lambda=0.01$, $M\in\left\{10,100\right\}$, and the number of local steps $K\in\left\{10,100,1000\right\}$. As predicted by our analysis, the performance of Minibatch SGD, Local SGD, and Local SGD-M degrades when the mixing time increases.
Dependent vs Independent Batches: The Markovian data streams detailed in Sec. 3.2 and investigated numerically in Sec. 5 focused on the scenario where the data stream $X_{m}$ for each client $m$ consisted of samples drawn from a time-homogeneous transition kernel $P_{m}$ and initial distribution $Q$. This scenario models situations where communication is performed every $K$ collected samples. On the other hand, samples may arrive during communication, which cannot be used in client updates. For sufficiently long communication periods, the consecutive batches of $K$ samples used for updates will be approximately independent. In the stationary case, the initial distribution $\mathcal{Q}=\pi$, and it can be shown that the analysis in Sec. 4 also applies to this scenario. We therefore carry out additional experiments to compare the performance of independent batches with batches drawn sequentially from the time-homogeneous data stream described in Sec. 3.2. We highlight that the data streams of different clients remain independent.
In Figure 7, we plot the gradient norm trajectory $\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}$ of Minibatch SGD, Local SGD, and Local SGD-M, with algorithm parameters $\gamma=0.01,\eta=0.001$, $\beta=0.1$, $100$ clients, the number of local steps $K\in\{10,100,1000\}$, and the mixing time for every client $\tau_{m}=100$. We consider the following scenarios:
  • Dependent Batches: All $TK$ samples observed by client $m$ are drawn sequentially from the Markov chain $X_{m}$ with transition kernel $P_{m}$ and stationary initial distribution $\mathcal{Q}=\pi$.
  • Independent Batches: At the beginning of each communication round, the samples $x_{t}^{m}$ are drawn from the Markov chain with transition kernel $P_{m}$, and initial state $x_{t}^{m,1}\sim\pi$ draw independently of $x_{t-1}^{m}$. As such, $x_{t}^{m}$ is independent of $x_{t-1}^{m}$, while the samples $(x_{t}^{m,k})_{k\in[K]}$ remain statistically dependent.
The experimental results in Figure 7 demonstrate that Minibatch SGD, Local SGD, and Local SGD-M, perform similarly for both dependent and independent batch sampling. This supports the validity of our analysis in both the dependent and independent batch scenarios.
(a) $M=10,K=10$
(b) $M=10,K=100$
(c) $M=10,K=1000$
(d) $M=100,K=10$
(e) $M=100,K=100$
(f) $M=100,K=1000$
Figure 3: Gradient norm as a function of the number of communication rounds for Local SGD, Minibatch SGD & Local SGD-M with $\tau_{m}=10$ for all $m\in[M]$, and for different number of clients and number of local steps.
(a) $M=10,K=10$
(b) $M=10,K=100$
(c) $M=10,K=1000$
(d) $M=100,K=10$
(e) $M=100,K=100$
(f) $M=100,K=1000$
Figure 4: Gradient norm as a function of the number of communication rounds for Local SGD, Minibatch SGD & Local SGD-M with $\tau_{m}=1000$ for all $m\in[M]$, and for different number of clients and number of local steps.

### v2

@@ D.3 Local SGD with Momentum
Hence, by using Young’s inequality:
@@
Here, in the last step, we use the triangle inequality and the sample-wise smoothness (Assumption 4.2).
@@
Using again Young’s inequality, we have:
@@
We start by using Young inequality $\left(a+b\right)^{2}\leq\left(1+\frac{1}{K-1}\right)a^{2}+Kb^{2}$. Then, for any $0\leq k\leq K-1$ we have:
[…]
@@
Hence, by using the triangle inequality, we have:
@@
Here, we use the sample-wise smoothness and the uniform bound on the gradient noise in the last step (Assumptions 4.24.3).
@@
With the condition on $\eta$, we have that $4\eta^{2}\beta^{2}L^{2}K\leq\frac{1}{K}<\frac{1}{K-1}$. Hence, by using the geometric sum formula, we have:
@@
In the proof of Local SGD, we have to use the heterogeneity assumption to bound the term $\frac{1}{M}\sum_{m}\mathbb{E}\left[\left\lVert\nabla F_{m}\left(w_{t}\right)\right\rVert^{2}\right]$. The use of momentum allows us to build a recursive bound for this term by the following lemma.
Lemma D.11.
@@
By using Young’s inequality and the smoothness of $F_{m}$, we have that:
@@
$\begin{split}\eta KL\leq&\left(\frac{1}{2\beta}\wedge\sqrt{\frac{C_{\infty}}{120\nu_{ps}MK\beta^{2}}}\wedge\sqrt{\frac{L\Delta_{0}}{360\beta^{3}TG_{0}}}\wedge\frac{1}{225}\frac{1}{\left(1-\beta\right)}\wedge\frac{1}{525}\frac{1}{\beta\gamma LT}\right),\\ \gamma\leq&\frac{1}{\sqrt{60}}\frac{\beta}{L}.\end{split}$ (125)
@@
$\begin{split}\frac{1}{T}\sum_{t=0}^{T-1}\Xi_{t}\leq\frac{7}{2}\frac{L\Delta_{0}}{\beta T}+31\frac{C_{\infty}\sigma^{2}}{\nu_{ps}MK}+\frac{1}{4}\sum_{t=0}^{T-1}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right].\end{split}$ (126)
@@
By Lemma D.11, we have that:
@@
Plugging (127) back into Lemma D.10 we have:
@@
By condition on the step size, we have that $\mathcal{B}\leq\frac{2}{15}$. Therefore, plugging (128) back into Lemma D.9, by noticing that $\mathbb{E}\left[\left\lVert\nabla F\left(w_{-1}\right)\right\rVert^{2}\right]=0$, gives:
@@
$\begin{split}120(\eta KL\beta)^{2}\sigma^{2}\leq&\frac{C_{\infty}\sigma^{2}}{\nu_{ps}MK},\\ 360\left(\eta KL\beta\right)^{2}G_{0}\leq&\frac{L\Delta_{0}}{\beta T}.\end{split}$ (131)
And by the law of total expectation, we have:
$\mathbb{E}\!\left[\left\lVert\bar{g}_{t}-\nabla F(w_{t})\right\rVert^{2}\right]$ $=\mathbb{E}\!\left[\mathbb{E}_{t}\!\left[\left\lVert\bar{g}_{t}-\nabla F(w_{t})\right\rVert^{2}\right]\right]$ (132)
[…]
Hence, by plugging (131)) and (133) back into (130) we have:
$\begin{split}\frac{1}{T}\sum_{t=0}^{T-1}\Xi_{t}\leq\frac{5}{4}\frac{\Xi_{-1}}{\beta T}+\frac{L\Delta_{0}}{\beta T}+\frac{31C_{\infty}\sigma^{2}}{\nu_{ps}MK}+\frac{1}{4}\sum_{t=0}^{T-1}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right].\end{split}$ (134)
Finally noticing that if we initialize $v_{0}=0$, $\Xi_{-1}=\mathbb{E}\left[\left\lVert\nabla F\left(w_{0}\right)\right\rVert^{2}\right]\leq 2L\Delta_{0}$ by the smoothness of $F$. Hence, we prove the lemma. ∎
@@
Theorem D.13 (Theorem 4.9 in the main paper).
@@
$\begin{split}\eta KL&\leq\mathcal{O}\left(\frac{1}{\beta}\wedge\sqrt{\frac{C_{\infty}}{\nu_{ps}MK\beta^{2}}}\wedge\sqrt{\frac{L\Delta_{0}}{\beta^{3}TG_{0}}}\wedge\frac{1}{\left(1-\beta\right)}\wedge\frac{1}{\beta\gamma LT}\right),\\ \gamma&\leq\mathcal{O}\left(\frac{\beta}{L}\right),\end{split}$ (135)
@@
$\begin{split}\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\mathcal{O}&\left(\frac{L\Delta_{0}}{\beta T}+\frac{C_{\infty}\sigma^{2}}{\nu_{ps}MK}\right).\end{split}$ (136)
@@
By rearranging the terms in Lemma D.8:
@@
Using the condition on $\gamma$ and replacing $\frac{1}{T}\sum_{t=0}^{T-1}\Xi_{t}$ by Lemma D.12 proves the theorem. ∎
From Theorem D.13, we can derive directly the communication and sample complexity of Local SGD-M.
Corollary D.14 (Corollary 4.10 in the main paper).
Under the conditions of Theorem D.13, the required number of local steps and communication rounds for Local SGD-M to achieve $\mathbb{E}[\|\nabla F(\hat{w}_{T})\|^{2}]\leq\epsilon^{2}$ are:
$K=\mathcal{O}\left(\frac{C_{\infty}\sigma^{2}}{\nu_{ps}M\epsilon^{2}}\right),~T=\mathcal{O}\left(\frac{L\Delta_{0}}{\beta\epsilon^{2}}\right)$ (138)
@@
The proof is a direct consequence of Theorem D.13 and follows the same steps in the proof of Corollary D.4. ∎
D.4 Uniformly Ergodic Markov Processes
A common assumption in the literature on Markov SGD is uniform geometric ergodicity (see Definition B.10).
Assumption D.15.
The system-level Markov chain is uniformly ergodic. In particular, for any $n\in\mathbb{N}$, we have:
$\sup_{x\in\Omega}\left\lVert P^{n}\left(x,.\right)-\pi\right\rVert_{\mathrm{TV}}\leq c\rho^{n},$ (139)
for some $c\leq\infty$ and $\rho<1$.
This assumption, which characterizes exponentially fast convergence to a unique stationary measure of Markov processes, is widely used in the literature of Stochastic Approximation with Markovian noise [3, 37, 15]. We highlight that in the main paper, we only assume that the clients’ Markov chains converge to a unique stationary measure, without any further assumption on convergence speed. In this section, we show that our analysis can be easily extended with this additional stronger assumption.
In particular, from [43, Proposition 3.4], we have that:
$\frac{1}{\nu_{ps}}\leq 2\tau,$ (140)
where $\tau$ is the mixing time of the system-level Markov chain (see Proposition B.11 for definition). According to Proposition B.12, we also have that:
$\tau\leq\lceil\log M+1\rceil\tau_{\max},$ (141)
where $\tau_{max}=\underset{m\in[M]}{\max}\tau_{m}$ the maximum mixing time of clients’ Markov processes.
Hence, a key implication of the uniform ergodicity assumption is a characterization of the pseudo-spectral gap in terms of the mixing time. That is, $1/\nu_{ps}$ is replaced by $\tau$ in the proof of Lemma D.1. As a consequence, for Minibatch SGD, Local SGD, and Local SGD-M, the new required number of local steps is:
$K=\tilde{\mathcal{O}}\left(\frac{C_{\infty}\tau_{\max}\sigma^{2}}{M\epsilon^{2}}\right).$ (142)
The communication complexity remains unchanged under the additional assumption of uniform ergodicity.
We note that uniform ergodicity can also be used to characterize $C_{\infty}$, as is discussed further in Section D.5.
D.5 Characterization of $C_{\infty}$
A key quantity in our analysis is $C_{\infty}$, which impacts the required number of local steps. In particular, recall
$C_{\infty}=\sup_{x\in\Omega}\operatorname*{ess\,sup}\left|\frac{\mathrm{d}P(x,\cdot)}{\mathrm{d}\pi}\right|.$ (143)
When $|\Omega|<\infty$, $C_{\infty}$ is given by
$C_{\infty}=\max_{x,y\in\Omega}\frac{P(x,y)}{\pi(y)},$ (144)
which can be explicitly computed for the case of two-state reversible Markov chains with $\Omega=\{0,1\}$ satisfying
$P(x,y)=\left\{\begin{array}[]{lr}p,&x\neq y\\ 1-p,&x=y.\end{array}\right$. (147)
In this case, $\pi(x)=\frac{1}{2},~x\in\{0,1\}$. It then follows that
$C_{\infty}=2\max\{1-p,p\},$ (148)
which is minimized for $p=\frac{1}{2}$, corresponding to independence between consecutive samples from the Markov chain.
For Markov chains with finite state space, where $|\Omega|<\infty$, $C_{\infty}$ can be bounded using the uniform ergodicity assumption in Assumption D.15. For an arbitrary $x\in\Omega$, under the assumption $P(x,\cdot)\ll\pi$, observe that
$\frac{1}{2}\sum_{y\in\Omega}|P(x,y)-\pi(y)|$ $=\frac{1}{2}\sum_{x\in\Omega}\pi(y)\left|\frac{P(x,y)}{\pi(y)}-1\right|$
$\geq\frac{1}{2}\pi_{\min}\left|\max_{y\in\Omega}\frac{P(x,y)}{\pi(y)}-1\right|$
$\geq\frac{1}{2}\pi_{\min}\left(\max_{y\in\Omega}\frac{P(x,y)}{\pi(y)}-1\right),$ (149)
where $\pi_{\min}=\min_{y\in\Omega}\pi(y)$. It then follows that for arbitrary $x\in\Omega$,
$\max_{y\in\Omega}\frac{P(x,y)}{\pi(y)}\leq\frac{2\|P(x,\cdot)-\pi\|_{\mathrm{TV}}}{\pi_{\min}}+1.$ (150)
Choosing $x=\arg\max_{u\in\Omega}\max_{y\in\Omega}\frac{P(u,y)}{\pi(y)}$, and applying the uniform ergodicity assumption in Assumption D.15, we then have
$C_{\infty}\leq\frac{2c\rho}{\pi_{\min}}+1.$ (151)
For finite Markov chains under the uniform ergodicity assumption, [43, Proposition 3.4] provides a relationship between the total variation distance and the pseudo spectral gap. Applying [43, Proposition 3.4] then yields
$C_{\infty}$ $\leq\frac{(1-\gamma_{ps})^{(1-1/\gamma_{ps})/2}\sqrt{\frac{1}{\pi_{\min}}-1}}{\pi_{\min}}+1$ (152)
$\leq\frac{e\sqrt{1/\pi_{\min}-1}}{\pi_{\min}}+1$ (153)
D.6 Alternative Bounds for Stochastic Gradient Errors in the Non-Stationary Case
In Lemma D.1, we obtained a bound on the stochastic gradient error in terms of $C_{\infty}$ by choosing the stationary distribution as a reference. In the following lemma, we obtain an alternative bound on the stochastic gradient error by choosing a different reference distribution. This bound can distinguish between stationary and non-stationary Markovian data processes.
Lemma D.16.
Let $\mu$ be the initial distribution of the system-level Markov chain defined in Section 3.2. We assume that for every $t\in\mathbb{N}$, and for every $x\in\Omega$, $P(x,.)$ is absolutely continuous with respect to $\mu P^{Kt}$, and we define:
$C_{q,t}=\underset{x\in\Omega}{\sup}\operatorname*{ess\,sup}\left\lvert\frac{dP(x,.)}{d(qP^{kt})}\right\rvert$$\infty$ (154)
If Assumptions 3.1 and 4.3 hold, then we have:
$\mathbb{E}_{t}\!\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F(w_{t})\right\rVert^{2}\right]\leq\frac{4C_{q,t}\sigma^{2}}{\nu_{ps}MK}+2C_{q,t}\sigma^{2}\left\lVert qP^{Kt}-\pi\right\rVert_{\mathrm{TV}}.$
Proof.
Let $\rho$ be a distribution such that $\mu P^{Kt}\ll\rho$ and $\pi\ll\rho$. We then have:
$\mathbb{E}_{t}\!\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F(w_{t})\right\rVert^{2}\right]$ (155)
$=\mathbb{E}_{qP^{kt}}\!\left[\frac{dP\left(x_{t-1}^{K-1},.\right)}{d(qP^{kt})}\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F(w_{t})\right\rVert^{2}\right]$ (156)
$\leq\left\lVert\frac{dP(x_{t-1}^{K-1},.)}{d\left(qP^{Kt}\right)}\right\rVert_{\pi,\infty}\mathbb{E}_{qP^{Kt}}\!\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F(w_{t})\right\rVert^{2}\right]\quad\textrm{(Holder's inequality)}$ (157)
[…]
$\leq C_{q,t}\mathbb{E}_{\pi}\!\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F(w_{t})\right\rVert^{2}\right]$ (159)
$+C_{q,t}\int_{\Omega}\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F(w_{t})\right\rVert^{2}\left\lvert\frac{\mathrm{d}\left(qP^{Kt}\right)}{\mathrm{d}\rho}-\frac{\mathrm{d}\pi}{\mathrm{d}\rho}\right\rvert\mathrm{d}\rho$ (160)
$\leq\frac{4C_{q,t}\sigma^{2}}{\nu_{ps}MK}+2C_{q,t}\sigma^{2}\left\lVert qP^{Kt}-\pi\right\rVert_{\mathrm{TV}}$ (161)
In the last step, we follow (78) to bound the squared error of the gradient estimator at the stationary measure and Assumption 4.3 together with the definition of the total variation norm (Appendix B,Definition B.7). ∎
Observe that in the stationary case, for every $t\in\mathbb{N}$ the additional term $\|qP^{Kt}-\pi\|_{\mathrm{TV}}$ is zero, and $C_{q,t}=C_{\infty}$, and we recover the bound in Lemma D.1. On the other hand, in the non-stationary case, while the total variation norm can be bound using [43, Proposition 3.4], obtaining a tighter bound than the one in Lemma D.1 is challenging due to the term $C_{q,t}$.
Appendix E Additional Experiments
E.1 Experimental Setting
Computing environment: All the experiments performed in this paper are run entirely on many different CPU clusters provided by the Grid’5000 testbed, with different types of CPU (e.g., Intel Xeon E5-2698, Intel Xeon E5-2620, Intel Xeon E5-2630). All the software packages and datasets used for experiments in this paper are open-sourced, with the exact version provided. For the main experiments, we use 10 different random seeds, and report the average together with the $95\%$ confidence interval. Further detailed instructions to run the experiments are provided in the supplementary material.
In Section 5, we evaluated the performance of Local SGD, Minibatch SGD, and Local SGD with Momentum on the Beijing Multi-Site Air-Quality dataset. In this section, we provide details on the data preprocessing and the learning problem.
The Beijing Multi-Site Air-Quality dataset consists of hourly measurements from 12 weather stations across China, collected between March 2013 and February 2017. In particular, the dataset consists of measurements of 6 pollutants (PM2.5, PM10, S02, CO2, CO, and O3) and 6 meteorological indicators (temperature, pressure, dew point, rainfall, wind speed, and wind direction). In total, there are more than 400,000 measurements of each variable.
Figure 3: Original time series for PM2.5 pollution, temperature, and SO2 pollution. Observe the periodicity in the data indicating seasonality.
Figure 4: Time series of PM2.5, temperature, and SO2 with seasonality removed, which is utilized (along with the other pollution and meteorological variables) for the experiments.
Figure 5: Auto-correlation of PM2.5, SO2, and temperature time series with seasonality removed. Observe that, particularly for the temperature, there are high levels of correlation at small lags, indicating dependence between adjacent samples.
Figure 6: Correlation matrix of the data at three stations.
As illustrated in Figure 3, the dataset has a strong seasonality effect. As we seek to experimentally validate our theoretical analysis, which requires convergence to a stationary distribution. As such, we estimate and remove the seasonality via the statsmodel python package [52]. We then drop the PM10 and WD (wind direction) features, fill in the missing values with their average values during the last month, and normalize the data. In Figure 4, we plot some measures after these steps of preprocessing.
The focus of our theoretical analysis is on Markovian data. Pollution and meteorological data are commonly modelled via Markov chains. To verify that a Markovian model is relevant to the dataset, it is necessary to verify that there is dependence between samples. To this end, in Fig. 5 we plot the autocorrelation of the PM2.5 concentration time series. Observe that while the autocorrelation decreases as the lag increases, which indicates mixing. On the other hand, the autocorrelation with a small lag remains high, indicating the presence of dependence and the need for non-independent and, in particular, Markovian models for the evolution of the PM2.5 concentration.
The learning problem we consider is linear regression with the PM2.5 measurements as the response, and the meteorological variables and the other pollution variables utilized as covariates. From the correlation matrices in Figure 6, we can see that there is a strong correlation between PM2.5 and CO, DEWP (dew point), and NO2, while other features are less relevant. Hence, we consider the following loss function:
$f\left(w;(x,y)\right)$ $=\left(w^{T}x-y\right)^{2}+\lambda r(w),$ (162)
$r(w)$ $=\frac{1}{2}\sum_{i=1}^{d=9}\frac{w_{i}^{2}}{1+w_{i}^{2}}.$ (163)
where $y\in\mathbb{R}$ is the PM2.5 measure, $x\in\mathbb{R}^{9}$ is the vector of other variables measured at the same hour, and $w\in\mathbb{R}^{d}$ is the model parameter. The non-convex regularizer encourages small weights for irrelevant features.
To allocate the data over different clients, for each client we randomly sample $n\in\{6,12\}$ consecutive months from the first $36$ months from all the sites with seasonality removal. The remaining $12$ months are reserved for testing.
E.2 Step Size Scaling with the Number of Local Steps
We remind here the upper bounds for Local SGD in Theorem 4.7:
$\begin{split}\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\mathcal{O}&\left(\frac{\Delta_{0}}{\eta KT}+\frac{C_{\infty}\sigma^{2}}{\nu_{ps}MK}\right.\left.+\frac{LK\eta\left(\theta^{2}+\sigma^{2}\right)}{\delta^{2}}\right).\end{split}$ (164)
In Section 5, we perform experiments with a fixed local step size $\eta$ to better illustrate the heterogeneity effect on Local SGD when $K$ increases. In this section, we perform the same experiments as in Figure 7, but with the local step size $\eta$ scaled as $b/K$, where $b\in\left\{0.1,0.01\right\}$. This replaces the fixed local step size $\eta=0.01$ in the original experiments. We report the results after $10$ different seeds together with the $95\%$ confidence interval in Table 2.
From the above upper bound, we note that, for moderate values of the initial suboptimality gap and heterogeneity:
  • If $K$ is large, then the bound is dominated by the heterogeneity term.
  • If $K$ is small, then the bound is dominated by the variance term.
Choosing the step size as $\eta\propto 1/K$ leaves the first and third error terms unchanged with $K$, while the second term decays with increasing $K$. Hence, when this second term dominates for small $K$, the gradient norm $\nabla F(w)||^{2}$ should decrease before flattening out at a positive plateau. The patterns shown in each row align with this qualitative prediction.
The comparison of the two rows also supports the fact that the "variance" term cannot be controlled by the step size, which was discussed in Appendix C. Indeed, for a fixed $K$, reducing the step size by a factor of $10$ does not improve the performance. On the contrary, the gradient norm appears to increase for smaller $\eta$. This can be explained by the first term in the bound, associated with the suboptimality gap, which scales inversely with the step size.
Table 2: Mean gradient norm of Local SGD
$\eta$ $K=10$ $K=50$ $K=100$
$0.1/K$ $00506\pm 0.00076 0.0$$0307\pm 0.00041 0.0$$00233\pm 0.00030$
$0.01/K$ $01248\pm 0.00279 0.0$$0921\pm 0.00210 0.0$$01190\pm 0.00245$
E.3 Additional Experiments on Synthetic Data
(a) $K=10$
(b) $K=100$
(c) $K=1000$